a biologist estimates that 50% of deer in the region carry a certain type of tick. for a sample of 300 deer selected at random, what is the chance that 155 or fewer deer have this tick?


1
Expert's answer
2021-05-11T07:06:27-0400

"p=0.5
,
\hat{p}=\dfrac{155}{300}=0.516,n=300"


Probability that 155 or fewer deer has a tick-

"=P(z\le\dfrac{\hat{p}-p}{\sqrt{\frac{p(1-p)}{n}}})\\[9pt]=P(z\le \dfrac{0.516-0.5}{\sqrt{\frac{0.5(1-0.5)}{300}}})\\[9pt]=P(z\le 0.554)\\[9pt]=0.69387"